# Adaptive Multi-Scale Decomposition Framework for Time Series Forecasting

### v1

@@ 5.2 Ablation Study and Analysis
Figure 4: (a) AMD guides the prediction by assigning greater weight to the dominant time pattern. (b) Memory usage (MB), training time (ms/iter), and MSE comparisons on the ETTh1 dataset. The input and predicted lengths are set to 512 and 96, respectively. $n$ denotes the number of predictors.
Observation 4: The balance of the TP-Selector is essential.
@@
We conduct experiments on the scaling rate of the load balancing loss, denoted by $\lambda_{1}$ in Eq. 11. As shown in Tab. 3, utilizing $\mathcal{L}_{\text{selector}}$ results in significantly improved performance, exceeding $11.2\%$ in terms of MSE compared to when $\lambda_{1}=0.0$. This underscores the crucial role of implementing load-balancing losses. Furthermore, the selector weights in the Fig. 4(a) do not tend to favor specific temporal patterns, addressing the load imbalance issue in sparse MoE structures.
Observation 5: AMD improves other TSF methods as a plugin.
Finally, we explore whether our proposed MoE-based method can yield an improvement in the performance of other TSF methods. We selected DLinear [13] and MTS-Mixers [16] as the baselines. After integrating the MDM and AMS modules, the predictive capabilities of all these models are enhanced as shown in Tab. 4, while maintaining the same computational resource requirements.
@@
6 Conclusion
In this paper, we propose the Adaptive Multi-Scale Decomposition (AMD) framework for time series forecasting to address the inherent complexity of time series data by decomposing it into multiple temporal patterns at various scales and adaptively aggregating these patterns. Comprehensive experiments demonstrate that AMD consistently achieves state-of-the-art performance in both long-term and short-term forecasting tasks across various datasets, showcasing superior efficiency and effectiveness. Looking ahead, we plan to further explore the integration of multi-scale information and expand the application of AMD as a backbone or plugin in other mainstream time series analysis tasks, such as imputation, classification, and anomaly detection.
@@ Appendix A Proof of Theorem 1
$\left|\frac{f_{1}(a)-f_{1}(b)}{a-b}\right|$ $=\frac{1}{d}\left|\frac{\sum_{j=ad-d+1}^{ad}f(j)-\sum_{j=bd-d+1}^{bd}f(j)}{a-b}\right|$ (12)
@@
$=$ $\left|\frac{f_{n-1}(a)-f_{n-1}(b)+\sum_{j=0}^{\big{[}\frac{L}{d^{n}}\big{]}}g_{n}(a)W_{n-1}(j,a)-\sum_{j=0}^{\big{[}\frac{L}{d^{i+1}}\big{]}}g_{n}(b)W_{n-1}(j,b)}{a-b}\right|$ (17)
$\leq$ $\left|\frac{f_{n-1}(a)-f_{n-1}(b)}{a-b}\right|+\left|\frac{\sum_{j=0}^{\big{[}\frac{L}{d^{n}}\big{]}}g_{n}(a)W_{n-1}(j,a)-\sum_{j=0}^{\big{[}\frac{L}{d^{n}}\big{]}}g_{n}(b)W_{n-1}(j,b)}{a-b}\right|$ (18)
@@ Appendix C Details of Experiments
C.1 Detailed Dataset Descriptions
@@ C.2 Baseline Models
(1) MTS-Mixers [16] is an MLP-based model utilizing two factorized modules to model the mapping between the input and the prediction sequence. The source code is available at https://github.com/plumprc/MTS-Mixers.
(2) DLinear [13] is an MLP-based model with just one linear layer, which unexpectedly outperforms Transformer-based models in long-term TSF. The source code is available at https://github.com/cure-lab/LTSF-Linear.
(3) TiDE [12] is a simple and effective MLP-based encoder-decoder model. The source code is available at https://github.com/thuml/Time-Series-Library.
(4) PatchTST [6] is a Transformer-based model utilizing patching and CI technique. It also enable effective pre-training and transfer learning across datasets. The source code is available at https://github.com/yuqinie98/PatchTST.
(5) Crossformer [9] is a Transformer-based model introducing the Dimension-Segment-Wise (DSW) embedding and Two-Stage Attention (TSA) layer to effectively capture cross-time and cross-dimension dependencies. The source code is available at https://github.com/Thinklab-SJTU/Crossformer.
(6) FEDformer [8] is a Transformer-based model proposing seasonal-trend decomposition and exploiting the sparsity of time series in the frequency domain. The source code is available at https://github.com/DAMO-DI-ML/ICML2022-FEDformer.
(7) TimesNet [21] is a CNN-based model with TimesBlock as a task-general backbone. It transforms 1D time series into 2D tensors to capture intraperiod and interperiod variations. The source code is available at https://github.com/thuml/TimesNet.
(8) MICN [20] is a CNN-based model combining local features and global correlations to capture the overall view of time series. The source code is available at https://github.com/wanghq21/MICN.
C.3 Metric Details
@@ C.4 Hyper-Parameter Selection and Implementation Details
Appendix D Extra Experimental Results
D.1 Full Results
Due to the space limitation of the main text, we place the full results of long-term forecasting in Tab. 7. To evaluate the generality of our proposed AMS, we conduct long-term forecasting on existing real-world multivariate benchmarks under different prediction lengths $S\in\{96,192,336,720\}$. The input sequence length is searched among $\{96,192,336,512,672,720\}$ to the best for a fairer comparison for all baselines.
Table 7: Full results of the long-term forecasting task. All the results are averaged from 4 different prediction lengths $T\in\{96,192,336,720\}$. To the best for a fairer comparison for all baselines, the input sequence length $L$ is searched among $\{96,192,336,512,672,720\}$.
Models AMD PatchTST Crossformer FEDformer TimesNet MICN DLinear MTS-Mixers
[6]
[9]
[8]
[21]
[20]
[13]
[16]
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Weather 96 0.145 0.198 0.149 0.198 0.153 0.217 0.238 0.314 0.172 0.220 0.161 0.226 0.152 0.237 0.156 0.206
192 0.190 0.240 0.194 0.241 0.197 0.269 0.275 0.329 0.219 0.261 0.220 0.283 0.220 0.282 0.199 0.248
336 0.242 0.280 0.245 0.282 0.252 0.311 0.339 0.377 0.280 0.306 0.275 0.328 0.265 0.319 0.249 0.291
720 0.315 0.332 0.314 0.334 0.318 0.363 0.389 0.409 0.365 0.359 0.311 0.356 0.323 0.362 0.336 0.343
avg 0.223 0.262 0.226 0.264 0.230 0.290 0.310 0.357 0.259 0.287 0.242 0.298 0.240 0.300 0.235 0.272
ETTh1 96 0.369 0.397 0.370 0.399 0.386 0.429 0.376 0.415 0.384 0.402 0.396 0.427 0.375 0.399 0.372 0.395
192 0.401 0.417 0.413 0.421 0.416 0.444 0.423 0.446 0.457 0.436 0.430 0.453 0.405 0.416 0.416 0.426
336 0.418 0.427 0.422 0.436 0.440 0.461 0.444 0.462 0.491 0.469 0.433 0.458 0.439 0.443 0.455 0.449
720 0.439 0.454 0.447 0.466 0.519 0.524 0.469 0.492 0.521 0.500 0.474 0.508 0.472 0.490 0.475 0.472
avg 0.407 0.424 0.413 0.431 0.441 0.465 0.428 0.454 0.458 0.450 0.433 0.462 0.423 0.437 0.430 0.436
ETTh2 96 0.274 0.337 0.274 0.337 0.628 0.563 0.332 0.374 0.340 0.374 0.289 0.357 0.289 0.353 0.307 0.354
192 0.351 0.383 0.339 0.379 0.703 0.624 0.407 0.446 0.402 0.414 0.409 0.438 0.383 0.418 0.374 0.399
336 0.375 0.411 0.329 0.380 0.827 0.675 0.400 0.447 0.452 0.452 0.417 0.452 0.448 0.465 0.398 0.432
720 0.402 0.438 0.379 0.422 1.181 0.840 0.412 0.469 0.462 0.468 0.426 0.473 0.605 0.551 0.463 0.465
avg 0.350 0.392 0.330 0.379 0.835 0.676 0.388 0.434 0.414 0.427 0.385 0.430 0.431 0.447 0.386 0.413
ETTm1 96 0.284 0.339 0.290 0.342 0.316 0.373 0.326 0.390 0.338 0.375 0.314 0.360 0.299 0.343 0.314 0.358
192 0.322 0.362 0.332 0.369 0.377 0.411 0.365 0.415 0.371 0.387 0.359 0.387 0.335 0.365 0.354 0.386
336 0.361 0.383 0.366 0.392 0.431 0.442 0.391 0.425 0.410 0.411 0.398 0.413 0.369 0.386 0.384 0.405
720 0.421 0.418 0.416 0.420 0.600 0.547 0.446 0.458 0.478 0.450 0.459 0.464 0.425 0.421 0.427 0.432
avg 0.347 0.375 0.351 0.381 0.431 0.443 0.382 0.422 0.400 0.406 0.383 0.406 0.357 0.379 0.370 0.395
ETTm2 96 0.167 0.258 0.166 0.256 0.421 0.461 0.180 0.271 0.187 0.267 0.178 0.273 0.167 0.260 0.177 0.259
192 0.221 0.295 0.223 0.296 0.503 0.519 0.252 0.318 0.249 0.309 0.245 0.316 0.224 0.303 0.241 0.303
336 0.270 0.327 0.274 0.329 0.611 0.580 0.324 0.364 0.321 0.351 0.295 0.350 0.281 0.342 0.297 0.338
720 0.356 0.382 0.362 0.385 0.996 0.750 0.410 0.420 0.497 0.403 0.389 0.409 0.397 0.421 0.396 0.398
avg 0.253 0.315 0.255 0.315 0.632 0.578 0.292 0.343 0.291 0.333 0.277 0.336 0.267 0.332 0.277 0.325
ECL 96 0.129 0.225 0.129 0.222 0.187 0.283 0.186 0.302 0.168 0.272 0.159 0.267 0.153 0.237 0.141 0.243
192 0.148 0.242 0.147 0.240 0.258 0.330 0.197 0.311 0.184 0.289 0.168 0.279 0.152 0.249 0.163 0.261
336 0.164 0.258 0.163 0.259 0.323 0.369 0.213 0.328 0.198 0.300 0.196 0.308 0.169 0.267 0.176 0.277
720 0.195 0.289 0.197 0.290 0.404 0.423 0.233 0.344 0.220 0.320 0.203 0.312 0.233 0.344 0.212 0.308
avg 0.159 0.254 0.159 0.253 0.293 0.351 0.207 0.321 0.192 0.295 0.182 0.297 0.177 0.274 0.173 0.272
Exchange 96 0.083 0.201 0.093 0.214 0.186 0.346 0.136 0.276 0.107 0.234 0.102 0.235 0.081 0.203 0.083 0.201
192 0.171 0.293 0.192 0.312 0.467 0.522 0.256 0.369 0.226 0.344 0.172 0.316 0.157 0.293 0.174 0.296
336 0.309 0.402 0.350 0.432 0.783 0.721 0.426 0.464 0.367 0.448 0.272 0.407 0.305 0.414 0.336 0.417
720 0.750 0.652 0.911 0.716 1.367 0.943 1.090 0.800 0.964 0.746 0.714 0.658 0.643 0.601 0.900 0.715
avg 0.328 0.387 0.387 0.419 0.701 0.633 0.478 0.478 0.416 0.443 0.315 0.404 0.297 0.378 0.373 0.407
Traffic 96 0.366 0.259 0.360 0.249 0.512 0.290 0.576 0.359 0.593 0.321 0.508 0.301 0.410 0.282 0.462 0.332
192 0.381 0.265 0.379 0.256 0.523 0.297 0.610 0.380 0.617 0.336 0.536 0.315 0.423 0.287 0.488 0.354
336 0.397 0.273 0.392 0.264 0.530 0.300 0.608 0.375 0.629 0.336 0.525 0.310 0.436 0.296 0.498 0.360
720 0.430 0.292 0.432 0.268 0.573 0.313 0.621 0.375 0.640 0.350 0.571 0.323 0.466 0.315 0.529 0.370
avg 0.393 0.272 0.391 0.264 0.535 0.300 0.604 0.372 0.620 0.336 0.535 0.312 0.434 0.295 0.494 0.354
Solar 96 0.175 0.228 0.224 0.278 0.181 0.240 0.201 0.304 0.219 0.314 0.188 0.252 0.289 0.337 0.284 0.325
192 0.188 0.235 0.253 0.298 0.196 0.252 0.237 0.337 0.231 0.322 0.215 0.280 0.319 0.397 0.307 0.362
336 0.201 0.246 0.273 0.306 0.216 0.243 0.254 0.362 0.246 0.337 0.222 0.267 0.352 0.415 0.333 0.384
720 0.203 0.249 0.272 0.308 0.220 0.256 0.280 0.397 0.280 0.363 0.226 0.264 0.356 0.412 0.335 0.383
avg 0.192 0.240 0.256 0.298 0.204 0.248 0.243 0.350 0.244 0.334 0.213 0.266 0.329 0.400 0.315 0.363
$1^{st}count$ 44 24 1 0 0 2 5 0
D.2 Robustness Evaluation
@@
96 0.1751$\pm$0.0003 0.2277$\pm$0.0003 0.1293$\pm$0.0002 0.2552$\pm$0.0002 0.3659$\pm$0.0004 0.2591$\pm$0.0003
192 0.1875$\pm$0.0004 0.2350$\pm$0.0003 0.1481$\pm$0.0003 0.2419$\pm$0.0001 0.3806$\pm$0.0003 0.2647$\pm$0.0003
336 0.2009$\pm$0.0003 0.2456$\pm$0.0002 0.1642$\pm$0.0003 0.2581$\pm$0.0002 0.3965$\pm$0.0004 0.2725$\pm$0.0003
720 0.2032$\pm$0.0003 0.2490$\pm$0.0003 0.1948$\pm$0.0002 0.2890$\pm$0.0002 0.4302$\pm$0.0004 0.2918$\pm$0.0002
D.3 Hyper-Parameter Sensitivity
@@
D.4 Extra Ablation Study
Ablation on MDM.
MDM is used to extract different temporal patterns. We conduct an ablation experiment by removing MDM. The results are as shown in the first row of Tab. 9 when the number of downsampling operations $n$ equals 00. It can be seen that removing the MDM module does not yield satisfactory prediction results.
Ablation on AMS.
AMS is based on a MoE structure to exploit information from every temporal pattern. We conducted an ablation experiment by replacing the TPSelector and predictors in AMS with a single predictor. The results, as shown in Tab. 10, demonstrate a significant decrease in prediction accuracy.
Table 10: Ablation on AMS results of only a single predictor.
Datasets
ETTh1 ETTm1 Weather Solar-Energy ECL
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
96 0.379 0.406 0.291 0.349 0.150 0.206 0.179 0.235 0.137 0.234
192 0.413 0.428 0.330 0.373 0.199 0.248 0.194 0.246 0.157 0.251
336 0.434 0.440 0.373 0.398 0.248 0.287 0.209 0.257 0.171 0.267
720 0.458 0.466 0.443 0.436 0.320 0.339 0.213 0.263 0.203 0.298
Dense MoE Strategy and Sparse MoE Strategy for AMD.
In theory, the information contained in each temporal pattern is useful and discarding any one of them would result in loss of information. Therefore, we adopt the dense MoE strategy, where dominant temporal patterns are given larger weights, while others are given smaller weights instead of being set to 00. Here, we conduct ablation experiments that compare the dense MoE with the sparse MoE to demonstrate this assertion. As shown in Tab. 11, compared to Dense MoE, Sparse MoE shows increased prediction errors. This observation highlights the consequence of omitting non-dominant temporal pattern information, which invariably leads to a degradation in performance.
Table 11: Ablation on Dense MoE Strategy and Sparse MoE Strategy, where (d) represents Dense MoE, and (s) represents Sparse MoE.
Datasets
ETTh1(d) ETTh1(s) Weather(d) Weather(s) ECL(d) ECL(s)
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
96 0.369 0.397 0.375 0.403 0.145 0.198 0.150 0.205 0.129 0.225 0.132 0.228
192 0.401 0.417 0.410 0.426 0.190 0.240 0.193 0.245 0.148 0.242 0.151 0.244
336 0.418 0.427 0.428 0.436 0.242 0.280 0.244 0.284 0.164 0.258 0.165 0.259
720 0.439 0.454 0.444 0.459 0.315 0.332 0.315 0.334 0.195 0.289 0.199 0.293
@@ Appendix E Discussions on Limitations and Future Improvement
(1) Normalization : Real-world time series always exhibit non-stationary behavior, where the data distribution changes over time. RevIn [41] is a normalization-and-denormalization method for TSF to effectively constrain non-stationary information (mean and variance) in the input layer and restore it in the output layer. The work has managed to improve the delineation of temporal dependencies while minimizing the influence of noise. In AMD, we also adopt the method of Revin. However, it struggles to adequately resolve the intricate distributional variations among the layers within deep networks, so we need to make further improvements to address this distributional shift.
@@
(2) Patching : Inspired by the utilization of local semantic context in computer vision (CV) and natural language processing (NLP), the technique of patching is introduced [6]. Since TS data exhibit locality, individual time steps lack the semantic meaning found in words within sentences. Therefore, extracting the local semantic context is crucial for understanding their connections. Additionally, this approach has the advantage of reducing the number of parameters. In AMD, we also develop a patching mechanism to extract recent history information. However, how to better exploit locality remains an issue that requires further research.
@@
(3) Frequency Domain Representation : TS data, characterized by their inherent complexity and dynamic nature, often contain information that is sparse and dispersed across the time domain. The frequency domain representation is proposed to promise a more compact and efficient representation of the inherent patterns. Related methods [8] still rely on feature engineering to detect the dominant period set. However, some overlooked periods or trend changes may represent significant events, resulting in information loss. In the future, we can explore adaptive temporal patterns mining in the frequency domain, thereby utilizing the Complex Frequency Linear module proposed by FITS [30] to mitigate the problem of large parameter sizes in MLP-based models when the look-back length is long.
@@
(4) Channel Independence (CI) and Channel Dependence (CD) : CI and CD represent a trade-off between capacity and robustness [42], with the CD method offering greater capacity but often lacking robustness when it comes to accurately predicting distributionally drifted TS. In contrast, the CI method sacrifices capacity in favor of more reliable predictions. PatchTST [6] achieves SOTA results using the CI approach. However, neglecting correlations between channels may lead to incomplete modelling. In AMD, we leverage the CI approach while integrating dependencies across different variables over time, thus exploiting cross-channel relationships and enhancing robustness. However, the trade-off between capacity and robustness is also a balance between generalization and specificity, which still requires further research.
@@
(5) Sequence Decomposition : The classical TS decomposition method [43] divides the complex temporal patterns into seasonal and trend components, thereby benefiting the forecasting process. In AMD, we go beyond the constraints of seasonal and trend-based time series decomposition. Instead, we develop an adaptive decomposition, mixing and forecasting module that fully exploits the information from different temporal patterns. However, the effectiveness of the adaptive decomposition module relies heavily on the availability and quality of historical data, which poses challenges in scenarios with limited or noisy data.

### v2

@@ Results.
Figure 5: Memory usage (MB), training time (ms/iter), and MSE comparisons on the Weather dataset. The input and predicted lengths are set to 512 and 96, respectively.
Dense MoE Strategy and Sparse MoE Strategy.
In theory, the information contained in each temporal pattern is useful and discarding any one of them would result in loss of information. Therefore, we adopt the dense MoE strategy, where dominant temporal patterns are given larger weights, while others are given smaller weights instead of being set to 0. We conduct ablation experiments as shown in Tab. 3. Compared to Dense MoE, Sparse MoE shows increased prediction errors. This observation highlights the consequence of omitting non-dominant temporal pattern information, which invariably leads to a degradation in performance.
Efficiency Analysis.
We thoroughly compare the training time and memory usage of various baselines on the Weather dataset, using the official model configurations and the same batch size. The results, shown in Fig. 5, indicate that the efficiency of AMD outperforms other Transformer-based models and MLP-based models with a relatively small number of parameters. Furthermore, as the number of predictors increases, the prediction error (MSE) gradually decreases, but the training time and memory usage increase significantly. To strike a balance, we set the number of predictors to 8 across all experiments.
Analysis on the balance of the TP-Selector.
@@
We conduct experiments on the scaling rate of the load balancing loss, denoted by $\lambda_{1}$ in Eq. 11. As shown in Tab. 3, utilizing $\mathcal{L}_{\text{selector}}$ results in significantly improved performance, exceeding $11.2\%$ in terms of MSE compared to when $\lambda_{1}=0.0$. This underscores the crucial role of implementing load-balancing losses. Furthermore, the selector weights in the Fig. 3 do not tend to favor specific temporal patterns, addressing the load imbalance issue in sparse MoE structures.
@@
Analysis on AMD as a plugin.
Finally, we explore whether our proposed MoE-based method can yield an improvement in the performance of other TSF methods. We selected DLinear (Zeng et al. 2023) and MTS-Mixers (Li et al. 2023b) as the baselines. After integrating the MDM and AMS modules, the predictive capabilities of all these models are enhanced as shown in Tab. 4, while maintaining the same computational resource requirements.
Conclusion
In this paper, we propose the Adaptive Multi-Scale Decomposition (AMD) framework for time series forecasting to address the inherent complexity of time series data by decomposing it into multiple temporal patterns at various scales and adaptively aggregating these patterns. Comprehensive experiments demonstrate that AMD consistently achieves state-of-the-art performance in both long-term and short-term forecasting tasks across various datasets, showcasing superior efficiency and effectiveness.
Acknowledgements
This work is supported by the National Key R&D Program of China (Grant no. 2022YFB4501704), the National Science Fundation of China (Grant no. 62472317), the Fundamental Research Funds for the Central Universities, the Shanghai Science and Technology Innovation Action Plan Project (Grant no. 22YS1400600 and 24692118300). This work is also supported in part by the National Natural Science Foundation of China, under Grant (62302309,), Shenzhen Science and Technology Program (JCYJ20220818101014030).
@@ Appendix A Proof of Theorem 1
$\left|\frac{f_{1}(a)-f_{1}(b)}{a-b}\right|$ (12)
$=$ $\frac{1}{d}\left|\frac{\sum_{j=ad-d+1}^{ad}f(j)-\sum_{j=bd-d+1}^{bd}f(j)}{a-b}\right|$ (13)
@@
$=$ $\left|\frac{1}{a-b}\right|\bigg{|}f_{n-1}(a)-f_{n-1}(b)+$ (18)
$\sum_{j=0}^{\big{[}\frac{L}{d^{n}}\big{]}}g_{n}(a)W_{n-1}(j,a)-\sum_{j=0}^{\big{[}\frac{L}{d^{i+1}}\big{]}}g_{n}(b)W_{n-1}(j,b)\bigg{|}$ (19)
$\leq$ $\left|\frac{f_{n-1}(a)-f_{n-1}(b)}{a-b}\right|+$ (20)
$\left|\frac{\sum_{j=0}^{\big{[}\frac{L}{d^{n}}\big{]}}g_{n}(a)W_{n-1}(j,a)-\sum_{j=0}^{\big{[}\frac{L}{d^{n}}\big{]}}g_{n}(b)W_{n-1}(j,b)}{a-b}\right|$ (21)
@@ Appendix C Details of Experiments
Detailed Dataset Descriptions
@@
(1) MTS-Mixers (Li et al. 2023b) is an MLP-based model utilizing two factorized modules to model the mapping between the input and the prediction sequence. The source code is available at https://github.com/plumprc/MTS-Mixers.
(2) DLinear (Zeng et al. 2023) is an MLP-based model with just one linear layer, which unexpectedly outperforms Transformer-based models in long-term TSF. The source code is available at https://github.com/cure-lab/LTSF-Linear.
(3) TiDE (Das et al. 2023) is a simple and effective MLP-based encoder-decoder model. The source code is available at https://github.com/thuml/Time-Series-Library.
(4) PatchTST (Nie et al. 2023) is a Transformer-based model utilizing patching and CI technique. It also enable effective pre-training and transfer learning across datasets. The source code is available at https://github.com/yuqinie98/PatchTST.
(5) iTransformer (Liu et al. 2024a) embeds each time series as variate tokens and is a fundamental backbone for time series forecasting. The source code is available at https://github.com/thuml/iTransformer.
(6) Crossformer (Zhang and Yan 2023) is a Transformer-based model introducing the Dimension-Segment-Wise (DSW) embedding and Two-Stage Attention (TSA) layer to effectively capture cross-time and cross-dimension dependencies. The source code is available at https://github.com/Thinklab-SJTU/Crossformer.
(7) FEDformer (Zhou et al. 2022b) is a Transformer-based model proposing seasonal-trend decomposition and exploiting the sparsity of time series in the frequency domain. The source code is available at https://github.com/DAMO-DI-ML/ICML2022-FEDformer.
(8) TimesNet (Wu et al. 2023) is a CNN-based model with TimesBlock as a task-general backbone. It transforms 1D time series into 2D tensors to capture intraperiod and interperiod variations. The source code is available at https://github.com/thuml/TimesNet.
(9) MICN (Wang et al. 2023) is a CNN-based model combining local features and global correlations to capture the overall view of time series. The source code is available at https://github.com/wanghq21/MICN.
Metric Details
@@
96 0.1751$\pm$0.0003 0.2277$\pm$0.0003 0.1293$\pm$0.0002 0.2242$\pm$0.0002 0.3659$\pm$0.0004 0.2591$\pm$0.0003
192 0.1865$\pm$0.0004 0.2350$\pm$0.0003 0.1471$\pm$0.0003 0.2379$\pm$0.0001 0.3806$\pm$0.0003 0.2647$\pm$0.0003
336 0.2003$\pm$0.0003 0.2456$\pm$0.0002 0.1602$\pm$0.0003 0.2531$\pm$0.0002 0.3965$\pm$0.0004 0.2695$\pm$0.0003
720 0.2032$\pm$0.0003 0.2490$\pm$0.0003 0.1928$\pm$0.0002 0.2860$\pm$0.0002 0.4292$\pm$0.0004 0.2918$\pm$0.0002
Hyper-Parameter Selection and Implementation Details
Appendix D Extra Experimental Results
Robustness Evaluation
Hyper-Parameter Sensitivity
@@ Appendix E Discussions on Limitations and Future Improvement
(1) Normalization : Real-world time series always exhibit non-stationary behavior, where the data distribution changes over time. RevIn (Kim et al. 2022b) is a normalization-and-denormalization method for TSF to effectively constrain non-stationary information (mean and variance) in the input layer and restore it in the output layer. The work has managed to improve the delineation of temporal dependencies while minimizing the influence of noise. In AMD, we also adopt the method of Revin. However, it struggles to adequately resolve the intricate distributional variations among the layers within deep networks, so we need to make further improvements to address this distributional shift.
@@
(2) Patching : Inspired by the utilization of local semantic context in computer vision (CV) and natural language processing (NLP), the technique of patching is introduced (Nie et al. 2023). Since TS data exhibit locality, individual time steps lack the semantic meaning found in words within sentences. Therefore, extracting the local semantic context is crucial for understanding their connections. Additionally, this approach has the advantage of reducing the number of parameters. In AMD, we also develop a patching mechanism to extract recent history information. However, how to better exploit locality remains an issue that requires further research.
@@
(3) Frequency Domain Representation : TS data, characterized by their inherent complexity and dynamic nature, often contain information that is sparse and dispersed across the time domain. The frequency domain representation is proposed to promise a more compact and efficient representation of the inherent patterns. Related methods (Zhou et al. 2022b) still rely on feature engineering to detect the dominant period set. However, some overlooked periods or trend changes may represent significant events, resulting in information loss. In the future, we can explore adaptive temporal patterns mining in the frequency domain, thereby utilizing the Complex Frequency Linear module proposed by FITS (Xu, Zeng, and Xu 2024) to mitigate the problem of large parameter sizes in MLP-based models when the look-back length is long.
@@
(4) Channel Independence (CI) and Channel Dependence (CD) : CI and CD represent a trade-off between capacity and robustness (Han, Ye, and Zhan 2023), with the CD method offering greater capacity but often lacking robustness when it comes to accurately predicting distributionally drifted TS. In contrast, the CI method sacrifices capacity in favor of more reliable predictions. PatchTST (Nie et al. 2023) achieves SOTA results using the CI approach. However, neglecting correlations between channels may lead to incomplete modelling. In AMD, we leverage the CI approach while integrating dependencies across different variables over time, thus exploiting cross-channel relationships and enhancing robustness. However, the trade-off between capacity and robustness is also a balance between generalization and specificity, which still requires further research.
@@
(5) Sequence Decomposition : The classical TS decomposition method (Holt 2004) divides the complex temporal patterns into seasonal and trend components, thereby benefiting the forecasting process. In AMD, we go beyond the constraints of seasonal and trend-based time series decomposition. Instead, we develop an adaptive decomposition, mixing and forecasting module that fully exploits the information from different temporal patterns. However, the effectiveness of the adaptive decomposition module relies heavily on the availability and quality of historical data, which poses challenges in scenarios with limited or noisy data.